OpenStudy (anonymous):

IF A(4,-5) and B(7,-9), find AB. @ganeshie8 @zepdrix @satellite73 @jim_thompson5910 @dumbcow @tkhunny @dan815 @abb0t @Compassionate @Destinymasha @SithsAndGiggles @nincompoop @paki @Whitemonsterbunny17 @Yahoo! @arabpride @perl @midhun.madhu1987

OpenStudy (anonymous):

that means the distance?

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

length rather?

OpenStudy (anonymous):

'the distance between points A and B'

OpenStudy (anonymous):

from 4 to 7 is 3 steps left from -5 to -9 is 4 steps down

OpenStudy (anonymous):

you got the mother of all right triangles there a 3 - 4 - 5 right triangle, making your distance 5

OpenStudy (anonymous):

Answer is 5 then?

OpenStudy (anonymous):

if you are a slave to formulas you can compute \[\sqrt{(7-4)^2+(-5+9)^2}=\sqrt{3^2+4^2}=\sqrt{9+16}=\sqrt{25}=5\] but that is way too much work

OpenStudy (anonymous):

and yes, it is five

OpenStudy (anonymous):

Yea lol

OpenStudy (anonymous):

I've got a few more similar questions, would it be ok if i post them?

OpenStudy (anonymous):

post in a new thread always better

OpenStudy (anonymous):

Ok